OpenStudy (anonymous):

What is 82.4% written as a decimal? A. 0.824 B. 8.24 C. 82.4 D. 824

OpenStudy (anonymous):

@jabez177

OpenStudy (danjs):

a percent is 'out of 100' put it over 100

OpenStudy (anonymous):

X% = X/100 so use this logic...

OpenStudy (anonymous):

So I multiply by 100

OpenStudy (rocketisawesome):

you move the decimal over two spaces to the left and get 0.824

OpenStudy (anonymous):

oh ok

OpenStudy (rocketisawesome):

when you want tp find answers to this move decimal point over to the left 2 times

OpenStudy (rocketisawesome):

*to*

OpenStudy (anonymous):

so its .824

OpenStudy (danjs):

\[\frac{ 82.4 }{ 100 } = \frac{ 824 }{ 1000 } = 0.824\] divide by 100, multiplied by 10/10, you have 824 thousandths, or 0.824 or simply move the decimal 2 places left

jabez177 (jabez177):

^ :D

OpenStudy (anonymous):

Thanks everyone sory i couldnt give out all the medals but if i could i would give everyone one

OpenStudy (rocketisawesome):

@sydneymmelin I told you the answer and gave you an explanation of how to do it before he did just to let you know. And do not forget to move the decimal point over two spaces to the left to get the answer.

OpenStudy (anonymous):

ok thanks @RocketIsAwesome